Multiple-Product Break-even, Break-Even Sales Revenue

Cherry Blossom Products Inc. produces and sells yoga-training products: how-to DVDs and a basic equipment set (blocks, strap, and small pillows). Last year, Cherry Blossom Products sold 13,500 DVDs and 4,500 equipment sets. Information on the two products is as follows:

DVDs

Equipment Sets

Price

$8

$25

Variable cost per unit

4

15

Total fixed cost is $99,750.

Suppose that in the coming year, the company plans to produce an extra-thick yoga mat for sale to health clubs. The company estimates that 9,000 mats can be sold at a price of $19 and a variable cost per unit of $11. Total fixed cost must be increased by $33,250 (making total fixed cost $133,000). Assume that anticipated sales of the other products, as well as their prices and variable costs, remain the same.

X

Part 1: Sales Mix Instructions and Part 2: Break-Even

REQUIRED:

1. What is the sales mix of DVDs, equipment sets, and yoga mats?
3:1:2 ?

2. Compute the break-even quantity of each product.

Break-even DVDs units
Break-even equipment sets units
Break-even yoga mats units

3a. Prepare an income statement for Cherry Blossom Products for the coming year.

Cherry Blossom Products Inc.
Income Statement
For the Coming Year
$
$
$

3b. What is the overall contribution margin ratio? Use the contribution margin ratio to compute overall break-even sales revenue. (Note: Round the contribution margin ratio to the nearest whole percent; round the break-even sales revenue to the nearest dollar.)

Overall contribution margin ratio %
Overall break-even sales revenue $

4. Compute the margin of safety for the coming year in sales dollars.
$
Answer #1
Part 1
Total of cherry's products =13,500+4,500+9,000 =27,000
Sales Mix for DVDs =13,500/27,000 =1/2 or 3/6
Sales Mix for Equipment Sets =4,500/27,000 =1/6
Sales Mix for DVDs =9,000/27,000 =1/3 or 2/6
So the Sales Mix ratio is 3:1:2
Part 2 DVDs Equipment Sets Yoga Mats
Sales Price per unit $                          8 $                             25 $                          19
Variable cost per unit $                          4 $                             15 $                          11
Contribution margin per unit $                          4 $                             10 $                            8
Fixed Cost allocated in ratio of 3:1:2 $                 66,500 $                      22,167 $                   44,333
Break-even units(Fixed Cost/Contribution margin per unit) 16625 2217 5542
Part 3.a Cherry Blossoms Products Inc.
Income Statement
For the Coming Year
DVDs Equipment Sets Yoga Mats Total
Sales $              1,08,000 $                   1,12,500 $                1,71,000 $        3,91,500
Variable costs $                 54,000 $                      67,500 $                   99,000 $        2,20,500
Contribution margin $                 54,000 $                      45,000 $                   72,000 $        1,71,000
Fixed Costs $                 66,500 $                      22,167 $                   44,333 $        1,33,000
Net Income $                -12,500 $                      22,833 $                   27,667 $           38,000
Part 3.a Overall Contribution margin ratio =$171,000 / $391,500 =44%
Break-even sales Revenue =$133,000 / 44% =$302,273
(Fixed Cost/Contribution margin ratio)
Part 4 Margin of Safety =$391,500 - $302,273 =$89,228